# How Does the Smoothness Approximation Method Facilitate Generalization for Federated Adversarial Learning?

### v1

@@ Appendix D Generalization Analysis of VFAL
$\leq$ $\dfrac{1}{m}\sum\limits_{j=1}^{m}e^{\frac{cL\sqrt{d}}{\gamma}K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+\dfrac{2}{m}\sum\limits_{j=1}^{m}e^{\frac{cL\sqrt{d}}{\gamma}K\eta_{t}}K\eta_{t}\mathbb{E}[v]+\dfrac{2K\eta_{t}e^{\frac{cL\sqrt{d}}{\gamma}K\eta_{t}}}{mn_{i}}\mathbb{E}\|\nabla\ell_{\gamma,i}(\theta_{i,k})\|$
@@
In the following denote $\ell(\theta,(\tilde{\mathbf{x}},y))=\ell(\mathbf{W},(\tilde{\mathbf{x}},y))\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{2}(f_{\mathbf{\theta}}(\tilde{\mathbf{x}})-y)^{2}$, where $\theta=\mathrm{vec}(\mathbf{W}),\theta\in\mathbb{R}^{d},\mathbf{W}\in\mathbb{R}^{r\times s},d=r\times s$
@@
Suppose Assumption hold, we have
@@
$\nabla_{\theta,\theta}^{2}\ell(\theta,\tilde{z})=\nabla f_{\theta}(\tilde{\mathbf{x}})\nabla f_{\theta}(\tilde{\mathbf{x}})^{\top}+\nabla^{2}f_{\theta}(\tilde{\mathbf{x}})(f_{\mathbf{W}}(\tilde{\mathbf{x}})-y)$
@@
$\nabla_{\theta}f_{\theta}(\tilde{\mathbf{x}})=\begin{pmatrix}u_{1}\tilde{\mathbf{x}}\phi^{\prime}\left(\langle(\theta)_{1},\tilde{\mathbf{x}}\rangle\right)\\ u_{2}\tilde{\mathbf{x}}\phi^{\prime}\left(\langle(\theta)_{2},\tilde{\mathbf{x}}\rangle\right)\\ \vdots\\ u_{s}\tilde{\mathbf{x}}\phi^{\prime}\left(\langle(\theta)_{s},\tilde{\mathbf{x}}\rangle\right)\end{pmatrix}\in\mathbb{R}^{rs}$
and $\nabla^{2}f_{\mathbf{\theta}}(\tilde{\mathbf{x}})\in\mathbb{R}^{rs\times rs}$ with
$\nabla^{2}_{\theta\theta}f_{\theta}(\tilde{\mathbf{x}})=\begin{pmatrix}u_{1}\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{\top}\phi^{\prime\prime}(\langle(\theta)_{1},\tilde{\mathbf{x}}\rangle)&0&0&\dots&0\\ 0&u_{2}\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{\top}\phi^{\prime\prime}(\langle(\theta)_{2},\tilde{\mathbf{x}}\rangle)&0&\dots&0\\ \vdots&\ddots&\ddots&\vdots&\vdots\\ 0&0&0&\dots&u_{s}\tilde{\mathbf{x}}\tilde{\mathbf{x}}^{\top}\phi^{\prime\prime}(\langle(\theta)_{s},\tilde{\mathbf{x}}\rangle)\end{pmatrix}$
Note that we then immediately have with $\mathbf{v}=(\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{s})\in\mathbb{R}^{rs}$ with $\mathbf{v}_{i}\in\mathbb{R}^{r}$
$\|\nabla^{2}_{\theta\theta}f_{\theta}(\tilde{\mathbf{x}})\|_{2}$ $=\max_{\mathbf{v}:\|\mathbf{v}\|_{2}\leq 1}\sum_{j=1}^{s}u_{j}\langle\mathbf{v}_{j},\tilde{\mathbf{x}}\rangle^{2}\phi^{\prime\prime}(\langle(\theta)_{j},\tilde{\mathbf{x}}\rangle)$
$\leq\frac{1}{\sqrt{s}}\|\tilde{\mathbf{x}}\|_{2}^{2}B_{\phi^{\prime\prime}}\max_{\mathbf{v}:\|\mathbf{v}\|_{2}\leq 1}\sum_{j=1}^{s}\|\mathbf{v}_{j}\|_{2}^{2}$
@@
We then see that the maximum Eigenvalue of the Hessian is upper bounded for any $\theta\in\mathbb{R}^{rs}$, that is
@@
$\leq 2(C_{x}^{2}+\rho^{2})B_{\phi^{\prime}}^{2}+\frac{2(C_{x}^{2}+\rho^{2})B_{\phi^{\prime\prime}}}{\sqrt{s}}(\sqrt{s}B_{\phi}+C_{y})$
@@
Suppose Assumption hold, we have
@@
$\nabla_{\theta\mathbf{x}}^{2}\ell(\theta,\tilde{z})=\nabla_{\theta}f_{\theta}(\tilde{\mathbf{x}})\nabla_{\mathbf{x}}f_{\theta}(\tilde{\mathbf{x}})^{\top}+\nabla^{2}_{\theta\mathbf{x}}f_{\theta}(\tilde{\mathbf{x}})(f_{\mathbf{W}}(\tilde{\mathbf{x}})-y)$
@@
$\nabla_{\mathbf{x}}f_{\theta}(\tilde{\mathbf{x}})=\begin{pmatrix}u_{1}(\theta)_{1}\phi^{\prime}\left(\langle(\theta)_{1},\tilde{\mathbf{x}}\rangle\right)\\ u_{2}(\theta)_{2}\phi^{\prime}\left(\langle(\theta)_{2},\tilde{\mathbf{x}}\rangle\right)\\ \vdots\\ u_{r}(\theta)_{r}\phi^{\prime}\left(\langle(\theta)_{r},\tilde{\mathbf{x}}\rangle\right)\end{pmatrix}\in\mathbb{R}^{rs}$
and $\nabla^{2}_{\theta\mathbf{x}}f_{\mathbf{\theta}}(\tilde{\mathbf{x}})\in\mathbb{R}^{s\times r\times r}$ with
$\nabla^{2}_{\theta\mathbf{x}}f_{\theta}(\tilde{\mathbf{x}})=\begin{pmatrix}\mu_{1}\left(\phi^{\prime\prime}(\langle(\theta)_{1},\tilde{\mathbf{x}}\rangle)\theta_{1}\tilde{\mathbf{x}}^{T}+\phi^{\prime}(\langle(\theta)_{1},\tilde{\mathbf{x}}\rangle)\mathbf{I}\right)\\ \mu_{2}\left(\phi^{\prime\prime}(\langle(\theta)_{2},\tilde{\mathbf{x}}\rangle)\theta_{2}\tilde{\mathbf{x}}^{T}+\phi^{\prime}(\langle(\theta)_{2},\tilde{\mathbf{x}}\rangle)\mathbf{I}\right)\\ \vdots\\ \mu_{s}\left(\phi^{\prime\prime}(\langle(\theta)_{s},\tilde{\mathbf{x}}\rangle)\theta_{s}\tilde{\mathbf{x}}^{T}+\phi^{\prime}(\langle(\theta)_{s},\tilde{\mathbf{x}}\rangle)\mathbf{I}\right)\end{pmatrix}.$
@@
$\|\nabla^{2}_{\theta\mathbf{x}}f_{\theta}(\tilde{\mathbf{x}})\|_{2}\leq$ $\sqrt{\sum_{k=1}^{s}\left(|u_{k}|\left(B_{\phi^{\prime}}+(C_{x}+\rho)B_{\phi^{\prime\prime}}\|\theta_{k}\|_{2}\right)\right)^{2}}$
$\leq$ $\left(B_{\phi^{\prime}}+(C_{x}+\rho)B_{\phi^{\prime\prime}}C_{\theta}\right)\sqrt{\sum_{k=1}^{s}u_{k}^{2}}$
$\leq$ $B_{\phi^{\prime}}+(C_{x}+\rho)C_{\theta}B_{\phi^{\prime\prime}}$
@@
$\|\nabla^{2}_{\theta\mathbf{x}}\ell(\theta,\tilde{z})\|_{2}\leq$ $\|\nabla_{\theta}f_{\theta}(\tilde{\mathbf{x}})\|_{2}\|\nabla_{\mathbf{x}}f_{\theta}(\tilde{\mathbf{x}})\|_{2}+\|\nabla^{2}_{\theta\mathbf{x}}f_{\theta}(\tilde{\mathbf{x}})\|_{2}|f_{\theta}(\tilde{\mathbf{x}})-y|$
$\leq$ $(C_{x}+\rho)C_{\theta}B_{\phi^{\prime}}^{2}+\big{(}B_{\phi^{\prime}}+(C_{x}+\rho)C_{\theta}B_{\phi^{\prime\prime}}\big{)}(\sqrt{s}B_{\phi}+C_{y})$
and therefor it’s $\zeta_{x}$-smooth where $\zeta_{x}=(C_{x}+\rho)C_{\theta}B_{\phi^{\prime}}^{2}+\big{(}B_{\phi^{\prime}}+(C_{x}+\rho)C_{\theta}B_{\phi^{\prime\prime}}\big{)}(\sqrt{s}B_{\phi}+C_{y})$ and is increasing with $\mathcal{O}(\rho\sqrt{s})$. ∎
@@
Under Assumption and given $i\in[m]$, for any $\theta$ we have
$\|\nabla R_{i}(\theta_{t})-\nabla R(\theta_{t})\|\leq(\zeta_{x}+6L)D_{i}$
@@
Suppose $\ell_{\gamma}(\theta)$ is a $\zeta_{\theta}$-gradient Lipschitz and non-convex function w.r.t $\theta$.
@@
Suppose $h(\theta)$ is a $\zeta_{\theta}$-gradient Lipschitz and non-convex function w.r.t $\theta$.
@@
Next we give the proof of Theorem 4
Theorem 4. Without loss of generality, we assume $4K\eta_{t}C_{0}\geq 1$ and $s\geq 16\eta_{t}^{2}T^{2}K^{2}(b^{\prime}H_{K})^{2}(1+2\eta_{t}\zeta_{\theta})^{2}$,where $b^{\prime}=C_{x}^{2}B_{\varphi^{\prime\prime}}(C_{x}B_{\varphi^{\prime}}+\sqrt{2C_{0}}),H_{K}=2\sqrt{K\eta_{t}C_{0}}$. Let the step size be chosen as $\eta_{t}\leq\frac{1}{6K\zeta_{\theta}}$. Under Assumptions 1-3, the generalization bound $\varepsilon_{gen}$ with the over-parameterized smoothness approximation satisfies:
@@
According to Lemma ,we further have
$\|\theta_{j,k+1}-\theta_{j,k+1}^{\prime}\|^{2}\leq\|\theta_{j,k}-\theta_{j,k}^{\prime}\|^{2}+\eta^{2}\|g_{j}(\theta_{j,k})-g_{j}(\theta_{j,k})\|^{2}+\eta_{t}^{2}(2\eta_{t}\zeta_{\theta}-3)\|g_{j}(\theta_{j,k})-g_{j}(\theta_{j,k})\|+2\eta\varsigma_{k}\|\theta_{j,k+1}-\theta_{j,k+1}^{\prime}\|^{2}$
@@
$\|\theta_{j,k+1}-\theta_{j,k+1}^{\prime}\|^{2}\leq\frac{1}{1-2\eta_{t}\varsigma_{t,k}}\|\theta_{j,k}-\theta_{j,k}^{\prime}\|^{2}$
@@
$\|\theta_{j,K}-\theta_{j,K}^{\prime}\|^{2}\leq$ $\prod_{k=0}^{K-1}\frac{1}{1-2\eta_{t}\varsigma_{t,k}}\|\theta_{t}-\theta_{t}^{\prime}\|^{2}$ (18)
where the second inequality can be found in Lei, Jin, and Ying (2022)(Theorem 7).
@@
$\leq$ $\frac{1}{1-2\eta_{t}\varsigma_{t,k}}\|\theta_{i,k}-\theta_{i,k}^{\prime}\|^{2}$
@@
$\leq$ $(1+p)\|\theta_{i,k+1}-\theta_{i,k+1}^{\prime}\|^{2}+2(1+\frac{1}{p})\eta_{t}^{2}\big{(}\|g(\theta_{i,k})\|^{2}+\|g(\theta_{i,k})^{\prime}\|^{2}\big{)}$
@@
$\mathbb{E}\|\theta_{i,k+1}-\theta_{i,k+1}^{\prime}\|^{2}\leq$ $(\frac{1}{1-2\eta_{t}\varsigma_{t,k}}+\frac{p}{n_{i}})\mathbb{E}\|\theta_{i,k}-\theta_{i,k}^{\prime}\|^{2}+\frac{4(1+1/p)\eta_{t}^{2}}{n_{i}}\mathbb{E}\|g(\theta_{i,k})\|^{2}$
@@
$\mathbb{E}\|\theta_{i,K}-\theta_{i,K}^{\prime}\|^{2}\leq$ $\prod_{k=0}^{K-1}(\frac{1}{1-2\eta_{t}\varsigma_{t,k}}+\frac{p}{n_{i}})\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|^{2}+\sum\limits_{k=0}^{K-1}\prod_{l=k+1}^{K-1}(\frac{1}{1-2\eta_{i,l}\varsigma_{t,l}}+\frac{p}{n_{i}})\dfrac{4(1+1/p)\eta_{t}^{2}}{n_{i}}\mathbb{E}\|g(\theta_{i,k})\|^{2}$ (19)
$\leq$ $\prod_{k=0}^{K-1}(\frac{1}{1-2\eta_{t}\varsigma_{t,k}}+\frac{p}{n_{i}})\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|^{2}+\sum\limits_{k=0}^{K-1}\dfrac{4e^{2}(1+1/p)\eta_{t}^{2}}{n_{i}}\mathbb{E}\|g(\theta_{i,k})\|^{2}$
By 18 and 19 we have
@@
$\leq$ $\frac{1}{m}\sum_{s=0}^{t-1}\sqrt{\sum_{k=0}^{K-1}\frac{4e^{4}(n_{i}+tK)}{n_{i}^{2}}}\mathbb{E}\|g(\theta_{i,k})\|$
@@
$\leq$ $\frac{1}{m}\sum_{t=0}^{T-1}\sqrt{\sum_{k=0}^{K-1}\frac{4e^{4}(n_{i}+tK)}{n_{i}^{2}}}(1+(1+\zeta_{\theta}K\eta_{t}))(\mathbb{E}\|\nabla R(\theta_{t})\|+(2\zeta_{x}+6L)D_{i}+\sigma)$
Under Theorem we complete the proof
@@
Appendix E SFAL Algorithms
Algorithm 4 SFAL of SSA
Output $\theta^{T}$ given by the server
1:  for $t=0,1...,T-1$ do
2:     $\theta^{t+1}_{i,0}=\theta_{t},\forall i=1,..,m$
3:     for $k=0,1,...,K-1$ do
5:        Accumulate the loss $\mathcal{L}_{loss}^{i.t}$ in the last local epoch
6:     end for
9:  end for
Algorithm 5 SFAL of RSA
Output $\theta^{T}$ given by the server
1:  for $t=0,1...,T-1$ do
2:     $\theta^{t+1}_{i,0}=\theta_{t},\forall i=1,..,m$
3:     Sample $\mathbf{u}^{t}\in\mathbb{R}^{d}$ uniformly from a unit sphere in $\mathbb{R}^{d}$ and send it to client $i$.
4:     for $k=0,1,...,K-1$ do
6:        Accumulate the loss $\mathcal{L}_{loss}^{i.t}$ in the last local epoch
7:     end for
10:  end for
Algorithm 6 SFAL of OPSA.
Output $\theta^{T}$ given by the server
1:  for $t=0,1...,T-1$ do
2:     $\theta^{t+1}_{i,0}=\theta_{t},\forall i=1,..,m$
3:     for $k=0,1,...,K-1$ do
5:        Accumulate the loss $\mathcal{L}_{loss}^{i.t}$ in the last local epoch
6:     end for
9:  end for
Appendix F Generalization Analysis of SFAL

### v2

@@ Appendix E Generalization Analyses of VFAL
$\leq$ $\dfrac{1}{m}\sum\limits_{j=1}^{m}e^{\frac{cL\sqrt{d}}{\gamma}K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+\dfrac{2}{m}\sum\limits_{j=1}^{m}e^{\frac{cL\sqrt{d}}{\gamma}K\eta_{t}}K\eta_{t}\mathbb{E}[v]+\dfrac{2K\eta_{t}e^{\frac{cL\sqrt{d}}{\gamma}K\eta_{t}}}{mn_{i}}\mathbb{E}\|\nabla\ell_{\gamma,i}(\theta_{i,k})\|.$
@@
In the following denote $\ell(\theta,(\tilde{x},y))=\ell(W,(\tilde{x},y))\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{2}(f_{W}(\tilde{x})-y)^{2}$. Unless stated otherwise, we work with vectorised quantities so $W\in\mathbb{R}^{ds}$. We also use notation $(W)_{\tau}$ so selected $\tau$-th block of size d, that is $(W)_{\tau}=[W_{(d-1)\tau+1},...,W_{d\tau}]^{\top}$.
@@
Suppose Assumption LABEL:Assumption_2-3 hold, we have
$\|\nabla^{2}_{WW}\ell(W,\tilde{z})\|_{2}\leq\zeta_{\theta},$
where $\zeta_{\theta}=2(C_{x}^{2}+\rho^{2})B_{\phi^{\prime}}^{2}+\frac{2(C_{x}^{2}+\rho^{2})B_{\phi^{\prime\prime}}}{\sqrt{s}}(\sqrt{s}B_{\phi}+C_{y}).$
@@
$\nabla_{W,W}^{2}\ell(W,\tilde{z})=\nabla f_{W}(\tilde{x})\nabla f_{W}(\tilde{x})^{\top}+\nabla^{2}f_{W}(\tilde{x})(f_{W}(\tilde{x})-y),$
@@
$\nabla_{\theta}f_{W}(\tilde{x})=\begin{pmatrix}\mu_{1}\tilde{x}\phi^{\prime}\left(\langle(W)_{1},\tilde{x}\rangle\right)\\ \mu_{2}\tilde{x}\phi^{\prime}\left(\langle(W)_{2},\tilde{x}\rangle\right)\\ \vdots\\ \mu_{s}\tilde{x}\phi^{\prime}\left(\langle(W)_{s},\tilde{x}\rangle\right)\end{pmatrix}\in\mathbb{R}^{ds},$
and $\nabla^{2}f_{W}(\tilde{x})\in\mathbb{R}^{ds\times ds}$ with
$\nabla^{2}_{WW}f_{W}(\tilde{x})=\begin{pmatrix}\mu_{1}\tilde{x}\tilde{x}^{\top}\phi^{\prime\prime}(\langle(W)_{1},\tilde{x}\rangle)&0&0&\dots&0\\ 0&\mu_{2}\tilde{x}\tilde{x}^{\top}\phi^{\prime\prime}(\langle(W)_{2},\tilde{x}\rangle)&0&\dots&0\\ \vdots&\ddots&\ddots&\vdots&\vdots\\ 0&0&0&\dots&\mu_{s}\tilde{x}\tilde{x}^{\top}\phi^{\prime\prime}(\langle(W)_{s},\tilde{x}\rangle)\end{pmatrix}.$
Note that we then immediately have with $v=(v_{1},v_{2},\ldots,v_{s})\in\mathbb{R}^{ds}$ with $v_{i}\in\mathbb{R}^{d}$
$\|\nabla^{2}_{WW}f_{W}(\tilde{x})\|_{2}$ $=\max_{v:\|v\|_{2}\leq 1}\sum_{\tau=1}^{s}\mu_{\tau}\langle v_{\tau},\tilde{x}\rangle^{2}\phi^{\prime\prime}(\langle(W)_{\tau},\tilde{x}\rangle)$
$\leq\frac{1}{\sqrt{s}}\|\tilde{x}\|_{2}^{2}B_{\phi^{\prime\prime}}\max_{v:\|v\|_{2}\leq 1}\sum_{\tau=1}^{s}\|v_{\tau}\|_{2}^{2}$
@@
We then see that the maximum Eigenvalue of the Hessian is upper bounded for any $W\in\mathbb{R}^{ds}$, that is
@@
$\leq 2(C_{x}^{2}+\rho^{2})B_{\phi^{\prime}}^{2}+\frac{2(C_{x}^{2}+\rho^{2})B_{\phi^{\prime\prime}}}{\sqrt{s}}(\sqrt{s}B_{\phi}+C_{y}),$
@@
Suppose Assumption LABEL:Assumption_2-3 hold, we have
$\|\nabla^{2}_{Wx}\ell(W,\tilde{z})\|_{2}\leq\zeta_{x},$
where $\zeta_{x}=(C_{x}+\rho)C_{W}B_{\phi^{\prime}}^{2}+\big{(}B_{\phi^{\prime}}+(C_{x}+\rho)C_{W}B_{\phi^{\prime\prime}}\big{)}(\sqrt{s}B_{\phi}+C_{y})$.
@@
$\nabla_{Wx}^{2}\ell(W,\tilde{z})=\nabla_{W}f_{W}(\tilde{x})\nabla_{x}f_{W}(\tilde{x})^{\top}+\nabla^{2}_{Wx}f_{W}(\tilde{x})(f_{W}(\tilde{x})-y),$
@@
$\nabla_{x}f_{W}(\tilde{x})=\begin{pmatrix}\mu_{1}(W)_{1}\phi^{\prime}\left(\langle(W)_{1},\tilde{x}\rangle\right)\\ \mu_{2}(W)_{2}\phi^{\prime}\left(\langle(W)_{2},\tilde{x}\rangle\right)\\ \vdots\\ \mu_{r}(W)_{r}\phi^{\prime}\left(\langle(W)_{d},\tilde{x}\rangle\right)\end{pmatrix}\in\mathbb{R}^{ds},$
and $\nabla^{2}_{Wx}f_{W}(\tilde{x})\in\mathbb{R}^{s\times d\times d}$ with
$\nabla^{2}_{Wx}f_{W}(\tilde{x})=\begin{pmatrix}\mu_{1}\left(\phi^{\prime\prime}(\langle(W)_{1},\tilde{x}\rangle)W_{1}\tilde{x}^{\top}+\phi^{\prime}(\langle(W)_{1},\tilde{x}\rangle)\mathbf{I}\right)\\ \mu_{2}\left(\phi^{\prime\prime}(\langle(W)_{2},\tilde{x}\rangle)W_{2}\tilde{x}^{\top}+\phi^{\prime}(\langle(W)_{2},\tilde{x}\rangle)\mathbf{I}\right)\\ \vdots\\ \mu_{s}\left(\phi^{\prime\prime}(\langle(W)_{s},\tilde{x}\rangle)W_{s}\tilde{x}^{\top}+\phi^{\prime}(\langle(W)_{s},\tilde{x}\rangle)\mathbf{I}\right)\end{pmatrix}.$
@@
$\|\nabla^{2}_{Wx}f_{W}(\tilde{x})\|_{2}\leq$ $\sqrt{\sum_{k=1}^{s}\left(|\mu_{k}|\left(B_{\phi^{\prime}}+(C_{x}+\rho)B_{\phi^{\prime\prime}}\|W_{k}\|_{2}\right)\right)^{2}}$
$\leq$ $\left(B_{\phi^{\prime}}+(C_{x}+\rho)B_{\phi^{\prime\prime}}C_{W}\right)\sqrt{\sum_{k=1}^{s}\mu_{k}^{2}}$
$\leq$ $B_{\phi^{\prime}}+(C_{x}+\rho)C_{W}B_{\phi^{\prime\prime}}.$
@@
$\|\nabla^{2}_{Wx}\ell(W,\tilde{z})\|_{2}\leq$ $\|\nabla_{W}f_{W}(\tilde{x})\|_{2}\|\nabla_{x}f_{W}(\tilde{x})\|_{2}+\|\nabla^{2}_{Wx}f_{W}(\tilde{x})\|_{2}|f_{W}(\tilde{x})-y|$
$\leq$ $(C_{x}+\rho)C_{W}B_{\phi^{\prime}}^{2}+\big{(}B_{\phi^{\prime}}+(C_{x}+\rho)C_{W}B_{\phi^{\prime\prime}}\big{)}(\sqrt{s}B_{\phi}+C_{y}),$
and therefor it’s $\zeta_{x}$-smooth where $\zeta_{x}=(C_{x}+\rho)C_{W}B_{\phi^{\prime}}^{2}+\big{(}B_{\phi^{\prime}}+(C_{x}+\rho)C_{W}B_{\phi^{\prime\prime}}\big{)}(\sqrt{s}B_{\phi}+C_{y})$ and is increasing with $\mathcal{O}(\rho\sqrt{s})$. ∎
@@
Under Assumption 1-3 and given $i\in[m]$, for any $\theta$ we have
$\|\nabla R_{i}(\theta_{t})-\nabla R(\theta_{t})\|\leq(\zeta_{x}+6L)D_{i},$
@@
Suppose $\ell_{\rho}(\theta)$ is a $\zeta_{\theta}$-gradient Lipschitz and non-convex function w.r.t $\theta$, we have
@@
Suppose $\ell_{\rho}(\theta)$ is a $\zeta_{\theta}$-gradient Lipschitz and non-convex function w.r.t $\theta$. we have
@@
Lemma 19.
(Almost Co-coercivity of the Gradient Operator(Lei, Jin, and Ying 2022)). Suppose Assumption LABEL:Assumption_2-3 hold, if $\eta\leq\frac{1}{2\zeta_{\theta}}$, then we have
$\langle\theta_{t}-\theta_{t}^{(i)},\ell(\theta_{t};z_{i})-\nabla\ell(\theta_{t}^{(i)};z_{i})\rangle$ $\geq 2\eta\Big{(}1-\frac{\eta\zeta_{\theta}}{2}\Big{)}\|\nabla\ell(\theta_{t};z_{i})-\nabla\ell(\theta_{t}^{(i)};z_{i})\|_{2}^{2}$
$-\varsigma_{t}\Big{\|}\theta_{t}-\theta_{t}^{(i)}-\eta\big{(}\nabla\ell(\theta_{t};z_{i})-\nabla\ell(\theta_{t}^{(i)};z_{i})\big{)}\Big{\|}_{2}^{2},$
where
$\varsigma_{t}=\frac{C_{x}^{2}B_{\phi^{\prime\prime}}}{\sqrt{m}}\Big{(}B_{\phi^{\prime}}C_{x}(1+2\eta\zeta_{\theta})\max\{\|\theta_{t}-\theta_{0}\|_{2},\|\theta_{t}^{(i)}-\theta_{0}\|_{2}\}+\sqrt{2C_{0}}\Big{)}.$
Next we give the proof of Theorem 4, for simplicity, we use $g(\theta)$ to denote strochastic gradient of loss function $\ell_{\rho}$
Theorem 4. Without loss of generality, we assume $4K\eta_{t}C_{0}\geq 1$ and $s\geq 16\eta_{t}^{2}T^{2}K^{2}(b^{\prime}H_{K})^{2}(1+2\eta_{t}\zeta_{\theta})^{2}$,where $b^{\prime}=C_{x}^{2}B_{\varphi^{\prime\prime}}(C_{x}B_{\varphi^{\prime}}+\sqrt{2C_{0}}),H_{K}=2\sqrt{K\eta_{t}C_{0}}$. Let the step size be chosen as $\eta_{t}\leq\frac{1}{6K\zeta_{\theta}}$. Under Assumptions 1-4, the generalization bound $\varepsilon_{gen}$ with the over-parameterized smoothness approximation satisfies:
@@
According to Lemma 19,we further have
$\|\theta_{j,k+1}-\theta_{j,k+1}^{\prime}\|^{2}\leq\|\theta_{j,k}-\theta_{j,k}^{\prime}\|^{2}+\eta^{2}_{t}\|g_{j}(\theta_{j,k})-g_{j}(\theta_{j,k})\|^{2}+\eta_{t}^{2}(2\eta_{t}\zeta_{\theta}-3)\|g_{j}(\theta_{j,k})-g_{j}(\theta_{j,k})\|+2\eta\varsigma_{k}\|\theta_{j,k+1}-\theta_{j,k+1}^{\prime}\|^{2}.$
@@
$\|\theta_{j,k+1}-\theta_{j,k+1}^{\prime}\|^{2}\leq\frac{1}{1-2\eta_{t}\varsigma_{t,k}}\|\theta_{j,k}-\theta_{j,k}^{\prime}\|^{2}.$
@@
$\|\theta_{j,K}-\theta_{j,K}^{\prime}\|^{2}\leq$ $\prod_{k=0}^{K-1}\frac{1}{1-2\eta_{t}\varsigma_{t,k}}\|\theta_{t}-\theta_{t}^{\prime}\|^{2}.$ (18)
@@
$\leq$ $\frac{1}{1-2\eta_{t}\varsigma_{t,k}}\|\theta_{i,k}-\theta_{i,k}^{\prime}\|^{2}.$
@@
$\leq$ $(1+p)\|\theta_{i,k+1}-\theta_{i,k+1}^{\prime}\|^{2}+2(1+\frac{1}{p})\eta_{t}^{2}\big{(}\|g(\theta_{i,k})\|^{2}+\|g(\theta_{i,k})^{\prime}\|^{2}\big{)}.$
@@
$\mathbb{E}\|\theta_{i,k+1}-\theta_{i,k+1}^{\prime}\|^{2}\leq$ $(\frac{1}{1-2\eta_{t}\varsigma_{t,k}}+\frac{p}{n_{i}})\mathbb{E}\|\theta_{i,k}-\theta_{i,k}^{\prime}\|^{2}+\frac{4(1+1/p)\eta_{t}^{2}}{n_{i}}\mathbb{E}\|g(\theta_{i,k})\|^{2}.$
@@
$\mathbb{E}\|\theta_{i,K}-\theta_{i,K}^{\prime}\|^{2}\leq$ $\prod_{k=0}^{K-1}(\frac{1}{1-2\eta_{t}\varsigma_{t,k}}+\frac{p}{n_{i}})\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|^{2}+\sum\limits_{k=0}^{K-1}\prod_{l=k+1}^{K-1}(\frac{1}{1-2\eta_{i,l}\varsigma_{t,l}}+\frac{p}{n_{i}})\dfrac{4(1+1/p)\eta_{t}^{2}}{n_{i}}\mathbb{E}\|g(\theta_{i,k})\|^{2}.$ (19)
Following the similar proof of Theorem 7 in (Lei, Jin, and Ying 2022), since $\|\theta_{i,k}-\theta_{i,0}\|\leq H_{K}$ and $\theta_{i,k}^{\prime}-\theta_{i,0}\|\leq H_{K}$, we know
$\varsigma_{t,k}\leq\frac{(1+2\eta_{t}\zeta_{\theta})b^{\prime}H_{K}}{\sqrt{m}}.$
Furthermore, $s\geq 16\eta_{t}^{2}T^{2}K^{2}(b^{\prime}H_{K})^{2}(1+2\eta_{t}\zeta_{\theta})^{2}$ implies $2\eta\varsigma_{t,k}\leq\frac{1}{tK+1}$ and therefore
$\prod_{l=k+1}^{K}\left(\frac{1}{1-2\eta_{t}\varsigma_{t,l}}+\frac{1}{K}\right)\leq\left(\frac{1}{1-1/(K+1)}+\frac{1}{K}\right)^{K}\leq\left(1+\frac{2}{K}\right)^{K}\leq e^{2}.$ (20)
$\prod_{l=s+1}^{t-1}\prod_{k=0}^{K-1}\left(\frac{1}{1-2\eta_{t}\varsigma_{l,k}}+\frac{1}{tK}\right)\leq\left(\frac{1}{1-1/(tK+1)}+\frac{1}{tK}\right)^{tK}\leq e^{2}.$ (21)
By 18,19,20 and 21, we have
@@
$\leq$ $\frac{1}{m}\sum_{s=0}^{t-1}\sqrt{\sum_{k=0}^{K-1}\frac{4e^{4}(n_{i}+tK)}{n_{i}^{2}}}\mathbb{E}\|g(\theta_{i,k})\|.$
@@
$\leq$ $\frac{1}{m}\sum_{t=0}^{T-1}\sqrt{\sum_{k=0}^{K-1}\frac{4e^{4}(n_{i}+tK)}{n_{i}^{2}}}(1+(1+\zeta_{\theta}K\eta_{t}))(\mathbb{E}\|\nabla R(\theta_{t})\|+(2\zeta_{x}+6L)D_{i}+\sigma).$
Combining Theorem 7 we get
@@
Discussion of the term including $K$ and $\sigma$. Similar to the discussion of the term including $\Delta$ in the main paper, the differences in the order of terms involving $K$ and $\sigma$ among three smoothness approximations also relate to their smoothing effects. For instance, in the RSA method, the term that includes both $K$ and $\sigma$ is optimized compared to SSA, owing to its derivation from a more flattened weight space.
Appendix F Generalization Analyses of SFAL
As we discuss in the main paper, SFAL only modifies the global aggregation method compared to VFAL, so we only give the proof of the theorem and Lemma related to global aggregation, and the rest is similar to the proof of VFAL.